OpenStudy (anonymous):

Find the value of: \[\large \lim_{n \rightarrow \infty } \left(\frac{1^{3}+2^{3}+...+n^{3}}{n^{4}}\right)\]

OpenStudy (anonymous):

\[e^{i \theta} = \cos(\theta)+isin(\theta)\] Can you work it out using this?

OpenStudy (anonymous):

Chances are you can.

OpenStudy (anonymous):

Yes you can, I'm trying to encourage him to work out the answer himself.

OpenStudy (anonymous):

\[\lim_{x \rightarrow infinity} \left(\frac{1^{3}+2^{3}+...+n^{3}}{n^{4}}\right) \]

OpenStudy (anonymous):

lol, sorry

OpenStudy (anonymous):

The answer isn't 0 lol.

OpenStudy (anonymous):

pardon?

OpenStudy (anonymous):

i never said that its 0..

OpenStudy (anonymous):

If that x under the limit was an n (which I assumed it was meant to be) then it would tend to 0.

OpenStudy (anonymous):

Anyway back to the question, are you able to work it out?

OpenStudy (anonymous):

i mean n tend to infinity

OpenStudy (anonymous):

Then that limit tends to 0, which isn't the right answer.

OpenStudy (anonymous):

its \[\large \lim_{n \rightarrow \infty } \left(\frac{1^{3}+2^{3}+...+n^{3}}{n^{4}}\right)\]

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

its not 0..

OpenStudy (anonymous):

how can it be 0 tho

OpenStudy (anonymous):

You have: \[\lim_{n \rightarrow \infty}(\frac{ \sum_{k=1}^{n}k^3 }{ k^4 })=\lim_{n \rightarrow \infty}( \sum_{k=1}^{n}\frac{ 1 }{ k } )=0\]

OpenStudy (anonymous):

Or to see it easier, if you divide each term on the top and bottom of the fraction by n^4 you will get lots of fractions all of which tend to 0 as n->infinity

OpenStudy (anonymous):

woops its\[\large \lim_{n \rightarrow \infty } \left(\frac{1^{3}+2^{3}+ 3^{3} ...+n^{3}}{n^{4}}\right)\] but i dont think its 0..

OpenStudy (anonymous):

its a Riemann sum change it to a simple integral

OpenStudy (anonymous):

Wait no I'm not, I just didn't type it in properly.

OpenStudy (anonymous):

the answer is 1/4

OpenStudy (anonymous):

hm..how do i change it into a simple integral?

OpenStudy (anonymous):

The top of the fraction, sum of k^3 from 1 to n, is 1/4 n^2(n+1)^2. So after cancelling you will get 1/4. Anyway, 1/4 is still not the answer to the original question so I don't see why we're talking about this...

OpenStudy (anonymous):

what math u are in Omn?

OpenStudy (anonymous):

is this from calculus?

OpenStudy (anonymous):

well..yes

OpenStudy (anonymous):

\[\large \lim_{n \rightarrow \infty } \left(\frac{1^{3}+2^{3}+ 3^{3} ...+n^{3}}{n^{4}}\right)=\lim_{n \rightarrow \infty } \left((\frac{1}{n})^3+(\frac{2}{n})^3+...+(\frac{n}{n})^3\right)\frac{1}{n}\\ \large=\int_{0}^{1}x^3 \text{d}x\]

OpenStudy (anonymous):

otherwise use the traxter method

OpenStudy (anonymous):

oohh..i got it thanks.

OpenStudy (anonymous):

:)